OpenStudy (anonymous):

Given cos 0 (with slash) = 1/2 and sin 0 (with slash) < 0, find sin 0 (with slash)Given cos 0 (with slash) = 1/2 and sin 0 (with slash) < 0, find sin 0 (with slash)@Mathematics

OpenStudy (stacey):

Do you know at what angle \[\cos \theta=1/2?\]

OpenStudy (anonymous):

Yes that's what I meant. And im not sure at what angle

OpenStudy (stacey):

Have you learned the unit circle? Or are you using triangles to solve these?

OpenStudy (anonymous):

We're using unit circle.

OpenStudy (stacey):

If you go to http://tutorial.math.lamar.edu/pdf/Trig_Cheat_Sheet.pdf , the third page has a unit circle.

OpenStudy (stacey):

The first number in each parenthesis is cos and the second is sin.

OpenStudy (stacey):

If you look for cos to be 1/2, that only occurs at 60 and 300 degrees.

OpenStudy (stacey):

Since according to your problem sin is less than zero, do you see what sin must equal?

OpenStudy (anonymous):

So sin would be -\[-\sqrt{3/2}\] ?

OpenStudy (anonymous):

Wrote that wrong sorry

OpenStudy (stacey):

Yes\[-\sqrt{3}/2\] That was a good attempt at using the equation editor. It takes some getting used to. :-)

OpenStudy (anonymous):

Lol yes it does. Thank you for all your help.

OpenStudy (stacey):

You're welcome.